The sum of two numbers is 23. The sum of the squares of the two numbers is 325. Find the numbers.
Answer by hello_me(3)

so x + y = 23
and x2 + y2 = 325
therefore using subsitution:
x=23-y
so
(23-y)2 + y2 = 325
so when we expand we get
529-46y+2y2 = 325
2y2-46y+204=102
by factorising we get
2(y-17)(y-6)=0
and therefore our numbers are 17 and 6